A certain concentration works best when the analyte’s concentration is approximately 10ppb.


B. If the analyte is present at 10% w/v, how would you prepare the sample for analysis


1
Expert's answer
2022-05-17T00:37:03-0400

c=10%=0.1

c1=10x10-9=10-8

c/c1=0.1/10-8=107

We have to dilute the sample 107 times

For example:

Place 1 ml of the sample in a 1 L flask and dilute to the mark with water. (solution 1). We dilute the sample 103 times.

Place 1 ml of the solution 1 in a 1 L flask and dilute to the mark with water. (solution 2). We dilute the sample 106 times.

Place 1 ml of the solution 2 in a 10 ml flask and dilute to the mark with water. We dilute the sample 107 times.